Henry's age is 10 years less than four times Paul's age. The sum of their ages is 100. Find their ages.
Answer by Alan3354(69443)

Henry's age is 10 years less than four times Paul's age. The sum of their ages is 100. Find their ages.
h = p+10
h+p = 100
sub for h:
p+10 + p = 100
2p+10 = 100
2p = 90
Paul is 45
Henry is 55